### v2

# A Banach space whose set of norm-attaining functionals is algebraically trivial

Abstract.
We construct a Banach space $\mathfrak{X}$ for which the set of norm-attaining functionals $\operatorname{NA}(\mathfrak{X},\mathbb{R})$ does not contain any non-trivial cone. Even more, given two linearly independent norm-attaining functionals on $\mathfrak{X}$, no other element of the segment between them attains its norm. Equivalently, the intersection of $\operatorname{NA}(\mathfrak{X},\mathbb{R})$ with the unit sphere of a two-dimensional subspace of $\mathfrak{X}^{*}$ contains, at most, four elements. In terms of proximinality, we show that for every closed subspace $Z$ of $\mathfrak{X}$ of codimension two, at most four element of the unit sphere of $\mathfrak{X}/Z$ has a representative of norm-one. We further relate this example with an open problem on norm-attaining operators.
Key words and phrases:
Banach space; bounded linear operator; Norm-attaining; proximinality; lineability
1991 Mathematics Subject Classification:
Primary 46B04, 46B20; Secondary 41A65, 46B03, 46B22, 46B87
1. Introduction
For a real Banach space $X$, we write $X^{*}$ for its topological dual space (endowed with the usual dual norm), and $B_{X}$, $S_{X}$ for its closed unit ball and its unit sphere, respectively. A (closed) subspace $M$ of $X$ is said to be proximinal if for every $x\in X\setminus M$, the set
(1) $P_{M}(x):=\{m\in M\colon\|x-m\|=\operatorname{dist}(x,M)\}$
is non-empty. This is equivalent to the fact that the quotient map $\pi_{M}\colon X\longrightarrow X/M$ satisfies that $S_{X/M}\subseteq\pi_{M}(S_{X})$. Clearly, every finite-dimensional subspace of $X$ is proximinal, and the same is true for subspaces which are reflexive. On the other hand, a hyperplane is proximinal if and only if it is the kernel of a norm-attaining functional. Recall that a functional $f\in X^{*}$ is norm-attaining if there exists $x\in S_{X}$ such that $\|f\|=|f(x)|$; we write $\operatorname{NA}(X,\mathbb{R})$ to denote the set of norm-attaining functionals on $X$. Observe that the Hanh-Banach theorem assures the existence of proximinal hyperplanes in every Banach space and, actually, the Bishop–Phelps theorem assures their abundance.
In the early 1970’s, Ivan Singer asked whether every (infinite dimensional) Banach space contains a proximinal subspace of codimension two (see [23, Problem 1], [24, Problem 2.1 in p. 14]). This question was answered negatively by the late Charles J. Read [21], who constructed an ingenious renorming $\mathcal{R}$ of $c_{0}$ without proximinal subspaces of codimension two, and so of any greater finite codimension. Singer’s problem is related to the following question by Gilles Godefroy of 2001 [7, Problem III] or [3, Question 2.26]: is it true that for every infinite-dimensional Banach space $X$, $\operatorname{NA}(X,\mathbb{R})$ contains two-dimensional subspaces? It is immediate that if $Y$ is a proximinal subspace of $X$ of finite-codimension, then $Y^{\perp}\subset\operatorname{NA}(X,\mathbb{R})$ (see [3, Lemma 2.2], for instance), but the reversed result is not true in general (see [11, Section 2] for instance). In [22, Theorem 4.2], Martin Rmoutil demonstrates that Read’s space $\mathcal{R}$ also gives a negative solution to Godefroy’s question. We refer the reader to [15] to find geometric properties of the space $\mathcal{R}$, to [16] to find further examples of spaces for which the set of norm-attaining functionals contains no two-dimensional subspace, and to [20] to find $n$-by-$n$ versions of these results. We do not know of other examples constructed in the literature which answer in the negative Singer’s question or Godefroy’s question.
Let us comment that both the space $\mathcal{R}$ and the main family of spaces constructed in [16] are not smooth. This implies that if $X$ is any of these spaces, there is an element $x\in S_{X}$ and two linearly independent elements $f_{1},f_{2}\in S_{X^{*}}$ such that $f_{1}(x)=f_{2}(x)=1$. Hence, $\operatorname{NA}(X,\mathbb{R})\cap S_{X^{*}}$ contains a non-trivial segment and, therefore, $\operatorname{NA}(X,\mathbb{R})$ contains non-trivial cones. There is a construction given in the proof of [6, Theorem 9.(4)] (see also [15, Lemma 11] and [10, Proposition 2.3]) which provides an equivalent smooth norm ||$\cdot$|| ⋅ | on a given separable Banach space $X$ such that $\operatorname{NA}((X,|\cdot|),\mathbb{R})=\operatorname{NA}(X,\mathbb{R})$. This allows to get a posteriori “Read spaces” which are smooth. Then, $\operatorname{NA}((X,|\cdot|),\mathbb{R})\cap S_{(X,|\cdot|),\mathbb{R})^{*}}$ does not contain non-trivial segments. However, $\operatorname{NA}((X,|\cdot|),\mathbb{R})$ contains non-trivial cones.
There is a somehow related open problem from the 1960’s of whether finite-rank operators can be always approximated by norm-attaining ones (see [13, Question 5], for instance). It is not even known whether there is a Banach space $X$ for which every non-zero norm-attaining operator from $X$ into the two dimensional Hilbert space $\ell_{2}^{2}$ is of rank-one. Here is the connection with the above discussion: it is shown in [17, Theorem 13.9] that the existence of non-trivial cones in $\operatorname{NA}(X,\mathbb{R})$ allows to construct a surjective norm-attaining operator from $X$ into $\ell_{2}^{2}$. It is natural to ask, as it was done in [17, Problem 13.11], whether $\operatorname{NA}(X,\mathbb{R})$ contains non-trivial cones for every Banach space $X$.
The aim of this paper is to construct a Banach space $\mathfrak{X}$ (which is actually a smooth renorming of $c_{0}$) for which $\operatorname{NA}(\mathfrak{X},\mathbb{R})$ contains no non-trivial cones. Moreover, the intersection of $\operatorname{NA}(\mathfrak{X},\mathbb{R})$ with any two-dimensional subspace $Z$ of $\mathfrak{X}^{*}$ is contained in the union of two one-dimensional subspaces (that is, the intersection of $S_{Z}$ with $\operatorname{NA}(\mathfrak{X},\mathbb{R})$ contains at most four points); equivalently, given two linearly independent elements in $\operatorname{NA}(\mathfrak{X},\mathbb{R})$, no element of the interior of the segment between them belongs to $\operatorname{NA}(\mathfrak{X},\mathbb{R})$ (all of these justify the name of algebraically trivial in the title of the paper). This is our Theorem 1. We also provide some remarks and comments on it:
  1. -
    The norm of $\mathfrak{X}$ can be get $\varepsilon$-equivalent to the given norm of $c_{0}$ for every $\varepsilon>0$ (Remark 2).
  2. -
    $\mathfrak{X}^{**}$ is strictly convex, hence $\mathfrak{X}^{*}$ is smooth and $\mathfrak{X}$ is also strictly convex (Proposition 3).
  3. -
    The construction doing in $c_{0}$ can be done in every weakly compactly generated (in particular, separable) Banach space which contains an isomorphic copy of $c_{0}$ and is (isomorphic to) a closed subspace of $\ell_{\infty}$ (Proposition 4).
We next deduce some proximinality results of the constructed space $\mathfrak{X}$ (see Proposition 6): given a closed subspace $M$ of $\mathfrak{X}$ of codimension two, the set $\{x\in B_{\mathfrak{X}}\colon\pi_{M}(x)\in S_{\mathfrak{X}/M}\}$ contains, at most, two linearly independent elements, hence $\pi_{M}(S_{X})\cap S_{X/M}$ contains, at most, two elements of $X/M$ and their opposites. Also, the set
$\{\pi_{M}(x)\colon x\in\mathfrak{X}\setminus M,\ P_{M}(x)\text{ is not empty}\}$
is contained in the union of two one-dimensional subspaces of $X/M$.
Finally, Subsection 2.3 contains an explanation of the relation of the constructed space $\mathfrak{X}$ and the problem of the denseness of norm-attaining finite-rank operators which was briefly explained above.
2. The example
This is the main result of the paper.
Theorem 1.
There is a Banach space $\mathfrak{X}$ such that for every two-dimensional subspace $Z$ of $\mathfrak{X}^{*}$, the intersection of $Z$ with $\operatorname{NA}(\mathfrak{X},\mathbb{R})$ is contained in the union of two one-dimensional subspaces (that is, the intersection of $S_{Z}$ with $\operatorname{NA}(\mathfrak{X},\mathbb{R})$ contains at most four points). Equivalently, given two linearly independent elements in $\operatorname{NA}(\mathfrak{X},\mathbb{R})$, no element of the interior of the segment between them belongs to $\operatorname{NA}(\mathfrak{X},\mathbb{R})$.
To prove the theorem, we will need a series of preliminary results, some of which may be of independent interest. We start constructing equivalent norms on $\ell_{1}$ which are smooth but still shares some properties of the usual duality between $\ell_{1}$ and $\ell_{\infty}$.
Lemma A.
Let $\Phi$ be a sequence of positive real numbers which belongs to $\ell_{1}$ with $\|\Phi\|_{1}<1$, and consider the operator $S_{\Phi}\colon\ell_{2}\longrightarrow\ell_{1}$ by $[S_{\Phi}(a)](n)=\Phi(n)a(n)$ for every $n\in\mathbb{N}$ and every $a\in\ell_{2}$. We define an equivalent norm $|\cdot|_{\Phi}$ on $\ell_{1}$ such that
$B_{(\ell_{1},|\cdot|_{\Phi})}=B_{\ell_{1}}+S_{\Phi}(B_{\ell_{2}}).$
Then:
  1. (a)
    for $f\in\ell_{\infty}$, $|f|_{\Phi}^{*}=\|f\|_{\infty}+\|S_{\Phi}^{*}(f)\|_{2}$;
  2. (b)
    for $f\in\ell_{\infty}$, $|f|_{\Phi}^{*}\geqslant\|f\|_{\infty}\geqslant(1-\|\Phi\|_{1})|f|_{\Phi}^{*}$;
  3. (c)
    for $x\in\ell_{1}$, $|x|_{\Phi}\leqslant\|x\|_{1}\leqslant(1+\|\Phi\|_{1})|x|_{\Phi}$;
  4. (d)
    $(\ell_{1},|\cdot|_{\Phi})^{*}$ is strictly convex, so $(\ell_{1},|\cdot|_{\Phi})$ is smooth;
  5. (e)
    if $x\in\ell_{1}$ and $f\in\ell_{\infty}$ satisfy that $|f|_{\Phi}^{*}=1$, $\langle f,x\rangle=|x|_{\Phi}$, and $|x(n)|>\Phi(n)|x|_{\Phi}$, it follows that $f(n)=\operatorname{sign}(x(n))\|f\|_{\infty}$.
Proof.
We follow the arguments in [6, Theorem 9.(4)] (given also at [15, Lemma 11]) and the more abstract version which appears in [16, Lemma 2.2]. First, the existence of the norm $|\cdot|_{\Phi}$ and the formula for the dual norm in (a) follows directly from [16, Lemma 2.2]. To get (b), just observe that $\|S_{\Phi}\|\leqslant\|\Phi\|_{1}$ and then
$\|S_{\Phi}^{*}(f)\|_{2}\leqslant\|S_{\Phi}\|\|f\|_{\infty}\leqslant\|S_{\Phi}\|\,|f|_{\Phi}^{*}.$
(c). $B_{\ell_{1}}\subset B_{(\ell_{1},|\cdot|_{\Phi})}\subset(1+\|S_{\Phi}\|)B_{\ell_{1}}\subset(1+\|\Phi\|_{1})B_{\ell_{1}}$.
(d) follows from the fact that $S_{\Phi}^{*}$ is one-to-one and the norm of $\ell_{2}$ is strictly convex using the formula given in (a).
(e). Suppose that $f\in S_{(\ell_{1},|\cdot|_{\Phi})^{*}}$ attains its norm at $x\in S_{(\ell_{1},|\cdot|_{\Phi})}$. Then, there is $x_{0}\in S_{\ell_{1}}$ and $a_{0}\in S_{\ell_{2}}$ such that $f$ attains its maximum on $B_{\ell_{1}}$ (equals to $\|f\|_{\infty}$) at $x_{0}$ and its maximum on $S_{\Phi}(B_{\ell_{2}})$ (equals to $\|S_{\Phi}^{*}(f)\|_{2}$) at $S_{\Phi}(a_{0})$, and $x=x_{0}+S_{\Phi}(a_{0})$. Now, for those $n\in\mathbb{N}$ such that $x_{0}(n)\neq 0$, it follows that $f(n)=\operatorname{sign}(x_{0}(n))\|f\|_{\infty}$. But if $x(n)>\Phi(n)$, then $x_{0}(n)\geqslant x(n)-\Phi(n)|a(n)|>\Phi(n)(1-|a(n)|)\geqslant 0$, hence $f(n)=\|f\|_{\infty}$. Analogously, if $x(n)<-\Phi(n)$, we get $f(n)=-\|f\|_{\infty}$. Finally, the result for $x$ with arbitrary norm follows by homogeneity. ∎
Next, we adapt Proposition 2.8 of [16] to get sequences which are “doubly infinitely dense” through an injective operator in any operator range. Recall that a linear subspace $Y$ of a Banach space $X$ is said to be an operator range if there is an infinite-dimensional Banach space $E$ and a bounded injective operator $T\colon E\longrightarrow X$ such that $T(E)=Y$. Equivalently, there is a complete norm on $Y$ which is stronger than the given norm of $Y$. We refer to the papers [4, 5, 12] and references therein for more information and background.
Lemma B.
Let $Y$ be a separable operator range. Then, there exists a norm-one injective operator $T\colon\ell_{1}(\mathbb{N}\times\mathbb{N})\longrightarrow Y$ such that for every $m\in\mathbb{N}$, the set
$\left\{\frac{T(e_{n,m})}{\|T(e_{n,m})\|}\colon n\in\mathbb{N}\right\}$
is dense in $S_{Y}$.
Proof.
Consider $\sigma\colon\mathbb{N}\times\mathbb{N}\longrightarrow\mathbb{N}$ bijective and the isometric isomorphism $\Psi\colon\ell_{1}(\mathbb{N}\times\mathbb{N})\longrightarrow\ell_{1}$ which carries $e_{n,m}\in\ell_{1}(\mathbb{N}\times\mathbb{N})$ to $e_{\sigma(n,m)}\in\ell_{1}$ for every $n,m\in\mathbb{N}$.
Now, take a sequence $\{w^{\prime}_{m}\colon m\in\mathbb{N}\}$ such that every element of $S_{Y}$ is an accumulation point of it and repeat the proof of Proposition 2.8 of [16] with the sequence $w_{\sigma(n,m)}=w^{\prime}_{m}$ for every $n,m\in\mathbb{N}$. ∎
We will also make use of the following two easy observations.
Observation C.
Let $\{Z_{m}\colon m\in\mathbb{N}\}$ be a family of smooth Banach spaces and let $Z=\left[\bigoplus_{m\in\mathbb{N}}Z_{m}\right]_{\ell_{1}}$ be its $\ell_{1}$ sum. If $z=(z_{m})\in Z$ and $f=(f_{m})\in Z^{*}=\left[\bigoplus_{m\in\mathbb{N}}Z_{m}^{*}\right]_{\ell_{\infty}}$ satisfy that $\|f\|_{\infty}=1$ and $\langle f,z\rangle=\|z\|_{1}$, then $f_{m}(z_{m})=\|z_{m}\|$ for every $m\in\mathbb{N}$, hence $\|f_{m}\|=1$ when $z_{m}\neq 0$. As a consequence, the norm of $Z$ is smooth at every element $z=(z_{m})\in X$ such that $z_{m}\neq 0$ for every $m\in\mathbb{N}$.
Observation D.
Let $Z$ be a smooth Banach space and let $z_{1}^{*},z_{2}^{*}\in\operatorname{NA}(Z,\mathbb{R})\cap S_{Z^{*}}$ be linearly independent. If for some $10<t<$ 1 we have $\|tz_{1}^{*}+(1-t)z_{2}^{*}\|=1$, then $tz_{1}^{*}+(1-t)z_{2}^{*}\notin\operatorname{NA}(Z,\mathbb{R})$.
We are now ready to prove the main result of the paper.
Proof of Theorem 1.
We divide the job in ten steps for clarity of the exposition.
Step 1.
[…]
$\operatorname{span}\operatorname{NA}((X,{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}),\mathbb{R})\cap Y=\{0\}.$
Step 2.
By Lemma B, there exists a norm-one operator $T\colon\ell_{1}(\mathbb{N}\times\mathbb{N})\longrightarrow Y\subset(X^{*},{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|})$ which is injective and satisfies that for every $m\in\mathbb{N}$, the set
(2) $\left\{\frac{T(e_{n,m})}{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|T(e_{n,m})\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}}\colon n\in\mathbb{N}\right\}$
is dense in $S_{Y}$ (hence, dense in $S_{(X,{\left|\kern-0.75346pt\left|\kern-0.75346pt\left|\cdot\right|\kern-0.75346pt\right|\kern-0.75346pt\right|})^{*}}$). Write $v^{*}_{n,m}:=T(e_{n,m})\in B_{Y}\subset B_{(X,{\left|\kern-0.75346pt\left|\kern-0.75346pt\left|\cdot\right|\kern-0.75346pt\right|\kern-0.75346pt\right|})^{*}}$ for every $n,m\in\mathbb{N}$.
Step 3.
Fix $m\in\mathbb{N}$. Let $|\cdot|_{m}$ be the equivalent norm on $\ell_{1}$ given by Lemma A for the sequence $\Phi_{m}(n)=\frac{1}{2^{m}}\frac{1}{2^{n}}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v_{n,m}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}$ for every $n\in\mathbb{N}$. Then:
  1. (i)
    $(\ell_{1},|\cdot|_{m})^{*}$ is strictly convex, so $(\ell_{1},|\cdot|_{m})$ is smooth;
  2. (ii)
    for $x\in\ell_{1}$, $|x|_{m}\leqslant\|x\|_{1}\leqslant\left(1+\frac{1}{2^{m}}\right)|x|_{m}$;
  3. (iii)
    for $f\in(\ell_{1},|\cdot|_{m})^{*}$, $|f(n)|\leqslant|f|_{m}^{*}$ for every $n\in\mathbb{N}$;
  4. (iv)
    if $x\in(\ell_{1},|\cdot|_{m})$, $f\in(\ell_{1},|\cdot|_{m})^{*}$, and $n\in\mathbb{N}$ satisfy that
    $|f|^{*}_{m}=1,\quad\langle f,x\rangle=|x|_{m},\quad|x(n)|>\frac{1}{2^{m}}\frac{1}{2^{n}}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v_{n,m}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}|x|_{m},$
    then
    $f(n)=\operatorname{sign}(x(n))\|f\|_{\infty}\quad\text{and}\quad\|f\|_{\infty}\in\left[1-\frac{1}{2^{m}},1\right].$
Step 4.
Define $R_{m}\colon(X,{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|})\longrightarrow(\ell_{1},|\cdot|_{m})$ by
$[R_{m}(x)](n)=\frac{m}{2^{m}}\frac{1}{2^{n}}v_{n,m}^{*}(x)\qquad\bigl{(}n\in\mathbb{N},\ x\in X\bigr{)}.$
Then:
  1. (a)
    $R_{m}$ is one-to-one since $\{v_{n,m}^{*}\colon n\in\mathbb{N}\}$ separates the points of $(X,{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|})$.
  2. (b)
    $|R_{m}(x)|_{m}\leqslant\|R_{m}(x)\|_{1}\leqslant\frac{m}{2^{m}}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|x\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}$ for every $x\in X$.
Step 5.
Define $R\colon(X,{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|})\longrightarrow W:=\left[\bigoplus\limits_{m\in\mathbb{N}}(\ell_{1},|\cdot|_{m})\right]_{\ell_{1}}$ by
$R(x)=\bigl{(}R_{m}(x)\bigr{)}_{m\in\mathbb{N}}\qquad(x\in X).$
Then, for $x\in X$,
$\|R(x)\|_{W}=\sum_{m=1}^{\infty}|R_{m}(x)|_{m}\leqslant\left(\sum_{m=1}^{\infty}\frac{m}{2^{m}}\right){\left|\kern-1.07639pt\left|\kern-1.07639pt\left|x\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}.$
By Step 4.(a), given $x\in X\setminus\{0\}$, all coordinates of $R(x)=\bigl{(}R_{m}(x)\bigr{)}_{m\in\mathbb{N}}$ are different from zero. This shows that $R$ is one-to-one and also that the norm of $W$ is smooth at $R(x)$ for every $x\in X\setminus\{0\}$ by Observation C.
Step 6.
Given $(w_{m}^{*})_{m\in\mathbb{N}}\in W^{*}\equiv\left[\bigoplus\limits_{m\in\mathbb{N}}(\ell_{1},|\cdot|_{m})^{*}\right]_{\ell_{\infty}}$, one has that
(3) $R^{*}\bigl{(}(w_{m}^{*})_{m\in\mathbb{N}}\bigr{)}=T\left(\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}w_{m}^{*}(n)\frac{m}{2^{m}}\frac{1}{2^{n}}e_{n,m}\right)\in Y\subset(X,{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|})^{*},$
where $T\colon\ell_{1}(\mathbb{N}\times\mathbb{N})\longrightarrow Y$ is the operator considered in Step 2; besides,
(4) $R^{*}\bigl{(}(w_{m}^{*})_{m\in\mathbb{N}}\bigr{)}=0\quad\Longrightarrow\quad w_{m}^{*}(n)=0\ \text{ for all $n,m\in\mathbb{N}$}.$
Indeed, Equation (3) follows routinely from the fact that the series
the series
$\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}w_{m}^{*}(n)\frac{m}{2^{m}}\frac{1}{2^{n}}e_{n,m}$
is absolutely sumable in $\ell_{1}$ as $|w_{m}^{*}(n)|\leqslant|w_{m}^{*}|_{m}^{*}\leqslant\left\|(w_{m}^{*})_{m\in\mathbb{N}}\right\|_{\infty}$ (by (iii) in Step 3). As $T$ is one-to-one and $\{e_{n,m}\colon n\in\mathbb{N},\,m\in\mathbb{N}\}$ is a basis of $\ell_{1}(\mathbb{N}\times\mathbb{N})$, Equation (4) follows immediately.
Step 7.
Define an equivalent norm $p$ on $X$ by
$p(x)={\left|\kern-1.07639pt\left|\kern-1.07639pt\left|x\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}+\|R(x)\|_{W}\qquad(x\in X)$
[…]
By Lemma 2.1.c in [16], if $x\in\mathfrak{X}$ and $x^{*}\in\mathfrak{X}^{*}$ satisfy that $p(x)=p(x^{*})=x^{*}(x)=1$, then there is $x^{*}_{0}\in\operatorname{NA}(X,{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|})$ with ${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|x_{0}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}=1$ and $w^{*}:=(w_{m}^{*})_{m\in\mathbb{N}}\in S_{W^{*}}$ such that
$x^{*}=x_{0}^{*}+R^{*}(w^{*})\quad\text{and}\qquad\langle(w_{m}^{*})_{m\in\mathbb{N}},R(x)\rangle=\|R(x)\|_{W}.$
Moreover, since $R_{m}(x)\neq 0$ for every $m\in\mathbb{N}$ (see Step 4.(a)), Observation C gives that
$|w_{m}^{*}|_{m}=1\quad\text{and}\quad w_{m}^{*}(R_{m}x)=|R_{m}x|_{m}\qquad(m\in\mathbb{N}).$
Step 8.
Pick $x^{*},z^{*}\in\operatorname{NA}(\mathfrak{X},\mathbb{R})$ with $p(x^{*})=p(z^{*})=1$ which are linearly independent and that attain their norms at $x,z\in S_{\mathfrak{X}}$, respectively. Hence, $x,z$ are linearly independent as the norm of $\mathfrak{X}$ is smooth. Suppose, for the sake of getting a contradiction, that $u^{*}:=tx^{*}+(1-t)z^{*}\in\operatorname{NA}(\mathfrak{X},\mathbb{R})$ for some $10<t<$ 1 and that it attains its norm at $u\in S_{\mathfrak{X}}$. Then, by Step 7, there are $x_{0}^{*},z_{0}^{*},u_{0}^{*}$ in $S_{(X,{\left|\kern-0.75346pt\left|\kern-0.75346pt\left|\cdot\right|\kern-0.75346pt\right|\kern-0.75346pt\right|})^{*}}\cap\operatorname{NA}(X,{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|})$, sequences $\Xi=(\xi_{m}^{*})_{m\in\mathbb{N}}$, $\Theta=(\zeta_{m}^{*})_{m\in\mathbb{N}}$, $\Upsilon=(u_{m}^{*})_{m\in\mathbb{N}}$ in $S_{W^{*}}$ satisfying
$x^{*}=x_{0}^{*}+R^{*}(\Xi),\quad z^{*}=z_{0}^{*}+R^{*}(\Theta),\quad\frac{tx^{*}+(1-t)z^{*}}{p(tx^{*}+(1-t)z^{*})}=u_{0}^{*}+R^{*}(\Upsilon),$
and
(5) $\xi_{m}^{*}(R_{m}x)=|R_{m}x|_{m},\quad\zeta_{m}^{*}(R_{m}z)=|R_{m}z|_{m},\quad u_{m}^{*}(R_{m}u)=|R_{m}u|_{m}\qquad(m\in\mathbb{N}).$
Therefore,
$tx_{0}^{*}+(1-t)z_{0}^{*}-p(tx^{*}+(1-t)z^{*})u_{0}^{*}=-R^{*}\bigl{(}t\,\Xi+(1-t)\Theta-p(tx^{*}+(1-t)z^{*})\Upsilon\bigr{)}.$
As the left-hand side belongs to $\operatorname{span}\operatorname{NA}(X,{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|})$, the right-hand side belongs to $Y$ by Step 6, and $\operatorname{span}\operatorname{NA}(X,{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|})\cap Y=\{0\}$ by Step 1, it follows that
$R^{*}\bigl{(}t\,\Xi+(1-t)\Theta-p(tx^{*}+(1-t)z^{*})\Upsilon\bigr{)}=0.$
Using now Equation (4), we get that
(6) $t\,\xi_{m}^{*}(n)+(1-t)\zeta^{*}_{m}(n)=p(tx^{*}+(1-t)z^{*})u_{m}^{*}(n)\quad\text{for all $n,m\in\mathbb{N}$}.$
Step 9.
As $x,z$ are linearly independent, we may find $\phi\in S_{(X,{\left|\kern-0.75346pt\left|\kern-0.75346pt\left|\cdot\right|\kern-0.75346pt\right|\kern-0.75346pt\right|})^{*}}$ and $\delta>0$ such that
$\phi(x)>\delta,\qquad\phi(z)>\delta.$
Indeed, just observe that $x$ does not belongs to the one-dimensional subspace $\mathbb{R}(x-z)$, and Hahn-Banach theorem gives $\phi\in S_{(X,{\left|\kern-0.75346pt\left|\kern-0.75346pt\left|\cdot\right|\kern-0.75346pt\right|\kern-0.75346pt\right|})^{*}}$ such that $\phi(x)=\operatorname{dist}(x,\mathbb{R}(x-z))$ and $\phi(x-y)=0$.
Step 10.
Consider $m\in\mathbb{N}$ such that
(7) $\frac{1}{m}<\delta\quad\text{and}\quad p(tx^{*}+(1-t)z^{*})<1-\frac{1}{2^{m}}$
(the second inequality is possible as Observation D gives that $p(tx^{*}+(1-t)z^{*})<1$ since $tx^{*}+(1-t)z^{*}\in\operatorname{NA}(\mathfrak{X},\mathbb{R})$). Now, as the set
$\left\{\frac{v^{*}_{n,m}}{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v^{*}_{n,m}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}}\colon n\in\mathbb{N}\right\}$
is dense in $S_{Y}$ by Step 2, hence dense in $S_{(X,{\left|\kern-0.75346pt\left|\kern-0.75346pt\left|\cdot\right|\kern-0.75346pt\right|\kern-0.75346pt\right|})^{*}}$, we may find $n\in\mathbb{N}$ such that
$v_{n,m}^{*}(x)>\delta{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v_{n,m}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\quad\text{and}\quad v_{n,m}^{*}(z)>\delta{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v_{n,m}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}.$
Then,
$[R_{m}x](n)$ $=\frac{m}{2^{m}}\frac{1}{2^{n}}v_{n,m}^{*}(x)>\frac{m}{2^{m}}\frac{1}{2^{n}}\delta{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v_{n,m}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}>\frac{1}{2^{m}}\frac{1}{2^{n}}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v_{n,m}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\geqslant\frac{1}{2^{m}}\frac{1}{2^{n}}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v_{n,m}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}|R_{m}x|_{m}.$
This together with Equation (5) imply by Step 3.(iv) that
$\xi_{m}^{*}(n)=\|\xi_{m}^{*}\|_{\infty}\in\left[1-\frac{1}{2^{m}},1\right].$
Analogously,
$\zeta_{m}^{*}(n)=\|\zeta_{m}^{*}\|_{\infty}\in\left[1-\frac{1}{2^{m}},1\right].$
By Equation (6), we have for the fixed $n,m\in\mathbb{N}$ that
$t\,\xi_{m}^{*}(n)+(1-t)\zeta^{*}_{m}(n)=p(tx^{*}+(1-t)z^{*})u_{m}^{*}(n).$
But the left-hand side belongs to the interval $\left[1-\frac{1}{2^{m}},1\right]$, while the right-hand side has absolute value strictly smaller than $1-\frac{1}{2^{m}}$ by Equation (7), a contradiction.∎
2.1. Remarks and further results on the construction
Let us gives some comments on the construction made above.
[…]
$m\|x\|\leqslant p(x)\leqslant M\|x\|\qquad(x\in X).$
Remark 2.
[…]
We next show some geometric properties of the space $\mathfrak{X}$.
Proposition 3.
The space $\mathfrak{X}$ constructed in the proof of Theorem 1 satisfies that $\mathfrak{X}^{**}$ is strictly convex, hence $\mathfrak{X}^{*}$ is smooth.
Proof.
We follow the lines of the proof of [16, Proposition 3.4]. To do so, we first observe that the bidual norm of the norm $p$ of $\mathfrak{X}$ given in Step 7 has the formula
$p(x^{**})=\|x^{**}\|_{X^{**}}+\|R^{**}(x^{**})\|_{W^{**}}\qquad(x^{**}\in\mathfrak{X}^{**})$
by [16, Lemma 2.1(b)]. Next, it is shown in Step 5 that the operator $R$ is one-to-one, but the same argument works for $R^{**}$ as for every $m\in\mathbb{N}$, the set $\{v_{n,m}^{*}\colon n\in\mathbb{N}\}$ separates the points of $(X,{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|\cdot\right|\kern-1.07639pt\right|\kern-1.07639pt\right|})^{**}$. Being $R$ a compact operator, we have that $R^{**}(\mathfrak{X}^{**})\subset J_{W}(R(X))$. Finally, if we show that $R(X)$ is strictly convex, then [16, Lemma 2.1(d)] will gives the desired strict convexity of $\mathfrak{X}^{**}$. Indeed, consider for the sake of contradiction, that there are $x,z\in S_{(X,{\left|\kern-0.75346pt\left|\kern-0.75346pt\left|\cdot\right|\kern-0.75346pt\right|\kern-0.75346pt\right|})}$ such that $R(x)$ and $R(z)$ are linearly independent and
(8) $\|R(x+z)\|_{W}=\|R(x)\|_{W}+\|R(x)\|_{W}.$
As $R$ is one-to-one, $x$ and $z$ are linearly independent and so, by the (weak-star) density of $S_{Y}$ in $S_{(X,{\left|\kern-0.75346pt\left|\kern-0.75346pt\left|\cdot\right|\kern-0.75346pt\right|\kern-0.75346pt\right|})^{*}}$, we may find $y^{*}\in S_{Y}$ and $m\in\mathbb{N}$ such that
$y^{*}(x)<-\frac{1}{m}<0<\frac{1}{m}<y^{*}(z).$
On the one hand, by the definition of the norm of $W$ in Step 5, it follows from Equation (8) that
$\bigl{|}R_{m}(x+z)\bigr{|}_{m}=|R_{m}(x)|_{m}+|R_{m}(z)|_{m}.$
Pick $f\in(\ell_{1},|\cdot|_{m})^{*}$ with $|f|_{m}^{*}=1$ satisfying that
$\langle f,R_{m}(x+z)\rangle=\bigl{|}R_{m}(x+z)\bigr{|}_{m}=|R_{m}(x)|_{m}+|R_{m}(z)|_{m}$
It then follows that
(9) $\langle f,R_{m}(x)\rangle=|R_{m}(x)|_{m}\qquad\text{and}\qquad\langle f,R_{m}(z)\rangle=|R_{m}(z)|_{m}.$
On the other hand, as the set $\left\{\dfrac{v^{*}_{n,m}}{{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v^{*}_{n,m}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}}\colon n\in\mathbb{N}\right\}$ is dense in $S_{Y}$ by Step 2, we may find $n\in\mathbb{N}$ such that
$v_{n,m}^{*}(x)<-\frac{1}{m}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v_{n,m}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\quad\text{and}\quad v_{n,m}^{*}(x)>\frac{1}{m}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v_{n,m}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}.$
Then,
$[R_{m}x](n)=\frac{m}{2^{m}}\frac{1}{2^{n}}v_{n,m}^{*}(x)<\frac{-1}{2^{m}}\frac{1}{2^{n}}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v_{n,m}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\leqslant\frac{-1}{2^{m}}\frac{1}{2^{n}}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v_{n,m}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\,|R_{m}x|_{m}$
and
$[R_{m}z](n)=\frac{m}{2^{m}}\frac{1}{2^{n}}v_{n,m}^{*}(z)>\frac{1}{2^{m}}\frac{1}{2^{n}}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v_{n,m}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\geqslant\frac{1}{2^{m}}\frac{1}{2^{n}}{\left|\kern-1.07639pt\left|\kern-1.07639pt\left|v_{n,m}^{*}\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}\,|R_{m}z|_{m}.$
This, together with Equation (9), imply by Step 3.(iv) that
$f(n)=-\|f\|_{\infty}<0\qquad\text{and}\qquad f(n)=\|f\|_{\infty}>0,$
a clear contradiction. ∎
The process done in Theorem 1 can be repeated in every weakly compactly generated Banach space $X$ which is isomorphic to a subspace of $\ell_{\infty}$ and contains an isomorphic copy of $c_{0}$.
Proposition 4.
Let $X$ be a weakly compactly generated (in particular, separable) Banach space which is isomorphic to a subspace of $\ell_{\infty}$ and contains an isomorphic copy of $c_{0}$. Then, there is an equivalent norm $p$ on $X$ such that for every two-dimensional subspace $Z$ of $(X,p)^{*}$, the intersection of $Z$ with $\operatorname{NA}((X,p),\mathbb{R})$ is contained in the union of two one-dimensional subspaces (that is, the intersection of $S_{Z}$ with $\operatorname{NA}((X,p),\mathbb{R})$ contains at most four points). Equivalently, given two linearly independent elements in $\operatorname{NA}((X,p),\mathbb{R})$, no element of the interior of the segment between them belongs to $\operatorname{NA}(\mathfrak{X},\mathbb{R})$.
Proof.
[…]
We do not know whether every non-reflexive Banach space can be renormed in such a way that the set of norm-attaining functionals in the new norm fails to contains non-trivial cones. However, having an equivalent norm with the properties of the one of $\mathfrak{X}$ in Theorem 1 is not possible for Banach spaces with the Radon-Nikodým property. Indeed, it is shown in [16, Proposition 5.6] that, in this case, the set of norm-attaining functionals contains the $\mathbb{Q}$-linear span of an $\mathbb{R}$-linearly independent infinite sequence. In particular, there are two-dimensional subspaces of the dual for which their intersection with the set of norm-attaining functionals is dense in the subspace.
The following remark on a result by Veselý of 2009 is pertinent.
Remark 5.
It was proved by L. Veselý in [25, Theorem 3.2] that for every Banach space $Z$, the set $\operatorname{NA}(Z,\mathbb{R})\cap S_{Z^{*}}$ is $\boldsymbol{c}$-dense in $S_{Z^{*}}$ (that is, every open subset of $S_{Z^{*}}$ contains at least continuum many elements of $\operatorname{NA}(Z,\mathbb{R})$). Moreover, if $Z$ admits an equivalent Fréchet differentiable norm (in particular, if $Z$ is isomorphic to $c_{0}$ as our space $\mathfrak{X}$ in Theorem 1), the set $\operatorname{NA}(Z,\mathbb{R})\cap S_{Z^{*}}$ is pathwise connected and locally pathwise connected. These results makes the existence of a space like $\mathfrak{X}$ even more surprising, if possible.
2.2. Some proximinality properties of the constructed space
It is clear that no two-codimensional subspace of $\mathfrak{X}$ is proximinal (as $\operatorname{NA}(\mathfrak{X},\mathbb{R})$ contains no two-dimensional subspaces). The next result shows that, actually, the situation is even more extreme.
Proposition 6.
Let $M$ be a two-codimensional closed subspace of $\mathfrak{X}$. Then:
  1. (a)
    The set
    $\bigl{\{}x\in B_{\mathfrak{X}}\colon\pi_{M}(x)\in S_{X/M}\bigr{\}}$
    contains, at most, two linearly independent elements.
  2. (b)
    There are two elements $x_{1},x_{2}\in B_{\mathfrak{X}}$ such that
    $\pi_{M}(B_{\mathfrak{X}})\cap S_{X/M}\subseteq\{x_{1}+M,x_{2}+M,-x_{1}+M,-x_{2}+M\}.$
  3. (c)
    There are two elements $x_{1},x_{2}\in\mathfrak{X}$ such that the (maybe empty) set of elements in the quotient $\mathfrak{X}/M$ given by
    $\bigl{\{}x+M\colon x\in\mathfrak{X}\setminus M,\ P_{M}(x)\text{ not empty}\bigr{\}}$
    is contained in $\mathbb{R}(x_{1}+M)\cup\mathbb{R}(x_{2}+M)$.
Proof.
(a). Take $x_{1},x_{2},x_{3}\in B_{\mathfrak{X}}$ satisfying that $\pi_{M}(x_{i})\in S_{X/M}$ for $i=1,2,3$. Pick $\phi_{1},\phi_{2},\phi_{3}\in(X/M)^{*}$ with $\phi_{i}(x_{i}+M)=\|\phi_{i}\|=1$ for $i=1,2,3$. Now, for $i=1,2,3$, writing $f_{i}:=\phi_{i}\circ\pi_{M}\in M^{\perp}\subset\mathfrak{X}^{*}$, we get that $\|f_{i}\|=f_{i}(x_{i})=1$ so, in particular, $f_{i}\in\operatorname{NA}(\mathfrak{X},\mathbb{R})$. As $M^{\perp}$ is a two-dimensional subspace of $\mathfrak{X}^{*}$, Theorem 1 gives that there are $i,j\in\{1,2,3\}$, $i\neq j$, such that $f_{i}=\pm f_{j}$. Being $\mathfrak{X}^{*}$ smooth by Proposition 3, we get that $x_{i}=\pm x_{j}$. Assertion (b) follows immediately from (a).
(c). Let $x_{1},x_{2}$ given by (b). Let $x\in\mathfrak{X}\setminus M$ such that $P_{M}(x)$ is not empty and take $m\in M$ such that $\|x-m\|=\operatorname{dist}(x,M)=\|x+M\|$. Then, $\dfrac{x-m}{\|x-m\|}\in S_{\mathfrak{X}}$ satisfies that for every $m^{\prime}\in M$,
$\left\|\dfrac{x-m}{\|x-m\|}-m^{\prime}\right\|=\frac{\bigl{\|}x-(m+\|x-m\|m^{\prime}\bigr{\|}}{\|x-m\|}\geqslant\frac{\operatorname{dist}(x,M)}{\|x-m\|}=1,$
hence $\pi_{M}\left(\dfrac{x-m}{\|x-m\|}\right)\in S_{\mathfrak{X}/M}$. It follows that
$\pi_{M}\left(\dfrac{x-m}{\|x-m\|}\right)\in\{\pm(x_{1}+M),\pm(x_{2}+M)\},$
that is, $\pi_{M}(x)\in\mathbb{R}(x_{1}+M)\cup\mathbb{R}(x_{2}+M)$. ∎
2.3. Relation with norm-attaining finite-rank operators
Given Banach spaces $X$, $Y$, we write $\mathcal{L}(X,Y)$ to denote the Banach space of all bounded linear operators from $X$ to $Y$ endowed with the operator norm. An operator $T\in\mathcal{L}(X,Y)$ is said to be norm-attaining (writing $T\in\operatorname{NA}(X,Y)$) if there is $x\in S_{X}$ such that $\|Tx\|=\|T\|$. Since the seminal paper by Lindenstrauss of 1963 [18], there has been an extensive study of when bounded linear operators can be approximated (in norm) by norm-attaining ones. This is the case if the domain is reflexive (Lindenstrauss), or even if it has the Radon-Nikodým property (Bourgain), or when the range is finite-dimensional and polyhedral or it is isometric to a subspace of $\ell_{\infty}$ containing the canonical copy of $c_{0}$ (Lindenstrauss). On the other hand, we may find operators which cannot be approximated by norm-attaining ones if either the domain is $c_{0}$ or $C[0,1]$ or $L_{1}[0,1]$ (Lindenstrauss) or the range is $\ell_{p}$ with $1\leqslant p<\infty$ (Gowers, Acosta). There are also compact operators which cannot be approximated by norm-attaining ones (Martín). The interested reader may find a lot of information in [1, 2, 14, 17, 19], and references therein.
Let us mention that it is still not known whether finite-rank operators can be always approximated by norm-attaining ones, even in the case when the range space is the two-dimensional Hilbert space $\ell_{2}^{2}$ (in the words of Johnson and Wolfe [13, Question 6], the latter is the most irritating open problem about norm-attaining operators). Actually, it is still unknown if there is a Banach space $X$ such that $\operatorname{NA}(X,\ell_{2}^{2})$ does not contain rank-two operators. Here is where the space $\mathfrak{X}$ comes into play in this setting. It is known that a finite rank operator $T\in\mathcal{L}(X,Y)$ belongs to $\operatorname{NA}(X,Y)$ whenever $[\ker T]^{\perp}\subseteq\operatorname{NA}(X,\mathbb{R})$ [17, Corollary 13.6]. Hence there are elements in $\operatorname{NA}(X,\ell_{2}^{2})$ of rank-two whenever $\operatorname{NA}(X,\mathbb{R})$ contains two-dimensional subspaces. Moreover, it is enough that $\operatorname{NA}(X,\mathbb{R})$ contains non-trivial cones to get that $\operatorname{NA}(X,Y)$ contains rank-two operators for every Banach space $Y$ of dimension greater than or equal to two [17, Corollary 13.10], and this applies to all constructed Read norms, as explained in the introduction, see [17, Example 13.13]. However, $\operatorname{NA}(\mathfrak{X},\mathbb{R})$ does not contain non-trivial cones, hence no known result implies that $\operatorname{NA}(\mathfrak{X},\ell_{2}^{2})$ contains rank-two operators.
Open Problem 7.
Does $\operatorname{NA}(\mathfrak{X},\ell_{2}^{2})$ contain rank-two operators?
Let us mention that [17, Theorem 13.24] contains a very general sufficient condition on $\operatorname{NA}(X,\mathbb{R})$ (which implies that it contains subspaces of arbitrary large dimension) to assure that every finite-rank operator whose domain is $X$ can be approximated by norm-attaining ones. Finally, we would like to send the interested reader attention to an open question related to the linear structure of the set of norm-attaining functionals [3, Question 2.25]: does there exists a non-reflexive Banach space $X$ for which $\operatorname{NA}(X^{*},\mathbb{R})$ is a vector space? It is known that in such a case, $X$ has to be weakly sequentially complete [3, Proposition 2.24]. More related open questions can be found in [9, §3]. To find applications of norm attainment, the reader is sent to [8].
Acknowledgments
The author has been supported by MICIU/AEI/10.13039/501100011033 and ERDF/EU through the grant PID2021-122126NB-C31, and by “Maria de Maeztu” Excellence Unit IMAG, reference CEX2020-001105-M funded by MICIU/AEI/10.13039/501100011033.
Thanks are also given to Audrey Fovelle, Gilles Godefroy, Vladimir Kadets, Ruben Medina, Abraham Rueda, and Dirk Werner for reading a first version of this manuscript and giving interesting remarks.